OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively.

OpenStudy (anonymous):

Need some helppp :(

OpenStudy (anonymous):

We know that the second term is: (a)(r) and the 5th term is: (a)(r^4) Also, 2304/(-36) is the 5th term divided by the 2nd term which is -64 That is also: [(a)(r^4)] / [(a)(r)] = r^3 So, r^3 = -64 and r = -4 So, the 2nd term, -36, is (a)(r) = (a)(-4), so "a" = 9 We now have "a" and "r" and the general expression for the nth term is: (9)(-4)^(n-1) And remember that exponentiation takes precedence over multiplication, so only the "-4" is being raised to the "n-1" power.

OpenStudy (anonymous):

All good now, @worne001 ?

OpenStudy (anonymous):

That problem is much more difficult than it appears, lol.. i guess I kind of understand..This is another exampled problem I have, kind of the same format.. Find an explicit rule for the nth term of the sequence. (is this the same process?) -5, -25, -125, -625, ...

OpenStudy (anonymous):

Okay, thank you so much ! :D

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @worne001 See you in the new post.